Calculate the radiation dosage (in rads) for an 79
Answer #1

number of disintegrations in 3Ci is

3.0 Ci ( 3.7 * 10^10 dintegrations/s/Ci )5 s = 5.55 * 10^11 disintegrations

Total energy of beta particle is

E_Beta =  5.55 * 10^11 disintegrations(2*10^-13 J/ disintegration) = 0.111 J

energy absorbed by 79 kg person is

E/m = 0.111 J/79 kg = 1.40 * 10^-3 gray ( 100 rad/ gray)= 0.140 rad